OpenStudy (anonymous):

(x + y)(x - y) ___________ (x + y) to the power of 2

jigglypuff314 (jigglypuff314):

(x+y) to the power of 2 = (x+y)*(x+y) yes? then a (x+y) from the top and bottom can cancel out so you are left with (x-y) over (x+y)

OpenStudy (anonymous):

so, ypu basically expand it and you get a (x-y) but i thought you would cancle out the bottom one and the answer would just be (x-y)

jigglypuff314 (jigglypuff314):

but (x+y) to the power of 2 = (x+y)*(x+y) meaning that there are two (x+y)'s multiplied by each other. and the top (x+y) only cancel's out one of the bottom ones.

OpenStudy (anonymous):

oh yeah i forgot that it was to the power of 2 thank you by the way i understand now

jigglypuff314 (jigglypuff314):

your welcome :)